How many grams of sulfosalicylic acid are needed to prepare 1 liter of 3% solution?


1
Expert's answer
2021-08-19T16:36:01-0400

"\%" "w\over v" "=" "g solute\over 100mL solution" "\times 100 \%"


"g solute=1mL \times \%"

"=1\times 3\over 100""=0.03 g"